OpenStudy (jmark):

If B=√3.a²/4 solve for 'a' and find the value of 'a' if B=16√3 ?

OpenStudy (anonymous):

a=8

OpenStudy (jmark):

@Addu5221 can you explain step by step procedure

OpenStudy (kamille):

\[16\sqrt{3}=\frac{ a ^{2}\sqrt{3} }{ 4 }|*4\] \[64\sqrt{3}=\sqrt{3}a ^{2}|:\sqrt{3}\ \[a ^{2}=64\] \[a=+/-8\]

OpenStudy (kamille):

Do you understand how I did it?

OpenStudy (anonymous):

its simple we have two values of B so they are equal so √3.a²/4=16√3 then a^2=4(16sqrt3)/sqrt3 a^2=64 then a=+8 or a=-8 since square of both - and + is +

OpenStudy (jmark):

now i understood thank you so much

OpenStudy (anonymous):

ur welcome